40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Oct 3 17:53:32 2017, awade, Craig, DailyProgress, FSS, Damage to the south interface board logic and debugging FSS north interface box power issues 
    Reply  Thu Oct 5 23:19:56 2017, awade, Craig, DailyProgress, FSS, Repairs to FSS interface boxes. 2017-10-05_16.13.12.jpg
       Reply  Fri Oct 6 21:18:19 2017, awade, DailyProgress, FSS, Repairs to FSS interface boxes. 
Message ID: 1946     Entry time: Fri Oct 6 21:18:19 2017     In reply to: 1944
Author: awade 
Type: DailyProgress 
Category: FSS 
Subject: Repairs to FSS interface boxes. 

The North and South path FSS field boxes have been connected and HV powered up.  

North locks and the cavity has very good absolute frequency stability with the vaccan temperature PID controls on. Its remained locked and well within the PZT actuator range for well over an hour. However, even with very low gains set, the PZT fast monitor is going into oscillations as it hits the rail.  I powered down the HV to disable the EOM and locked just with PZT and it was possible to turn the gain down and get a well behaved locking error signal.  My suspicion is that the EOM is fighting the PZT and there is some kind of sign error between the two. We can only flip the sign of the fast path. Maybe it is possible that PZT is correct and EOM is the wrong sign. We have a phase adjuster box, this is something to try: flip the overall sign of the error signal and then the field box switch.

We need to take a transfer function to see what is really going on, but this will only be possible if the error signal isn't entering into wild oscillations. 

The south path has some issues.  It is very likely that the alignment was bumped as I can't find a TEM00 mode anywhere. This needs to be aligned over the weekend.


In the previous setup I had put LP filters on the output of the FSS fast monitor before feeding into the acromag slow channels.  I'm now not sure if this is the right thing to do, re ADC noise.  Previously the slow channels fed into differential voltage channels on the acromag X1221 cards had not been grounded relative to the card ground return.  This had lead to a bunch of junk noise and pickup. The acromag manual recommends making a link to return from minus or plus terminal of each channel.  I suspect putting the LP filters in effectively filtered this with the capacitor of the LP across those terminals.  Still, without LP filtering the PID slow controls for the laser temperature bump the PDH lock off lock whenever the loop starts to ring. With the LP filters in, it just rides over the oscillations (giving just the average I guess) and lets the PID pull PZT to the center of its range.  Not sure if it is the 'right' thing to do to just put the LP filters back in?

The south FastMon in is properly linked to ground on the negative terminal, but it looks the North FastMon is still floating.

ELOG V3.1.3-